# Graph Max Shift: A Hill-Climbing Method for Graph Clustering

### v1

@@ Abstract
We present a method for graph clustering that is analogous with gradient ascent methods previously proposed for clustering points in space. We show that, when applied to a random geometric graph with data iid from some density with Morse regularity, the method is asymptotically consistent. Here, consistency is understood with respect to a density-level clustering defined by the partition of the support of the density induced by the basins of attraction of the density modes.
@@ 1 Introduction
A hill-climbing algorithm is typically understood as an algorithm that makes ‘local’ moves. In a sense, this class of procedures is the discrete analog of the class of gradient-based and higher-order methods in continuous optimization. Such algorithms have been proposed in the context of graph partitioning, sometimes as a refinement step, where the objective function is typically a notion of cut and local moves often take the form of swapping vertices in order to improve the value of the objective function. More specifically, consider an undirected graph consisting of $n$ nodes, which we take to be $[n]:=\{1,...,n\}$ without loss of generality, and adjacency matrix $A=(a_{ij})$, so that $a_{ij}=1$ when $i$ and $j$ are connected in the graph, and $a_{ij}=0$ otherwise. Consider the simplest variant of graph partitioning, where assuming $n$ is even, the goal is to bisect the node set into two subsets, $I\subset[n]$ and $I^{\mathsf{c}}$, in order to minimize the cut: ${\rm cut}(I)=\sum_{i\in I}\sum_{j\in I^{\mathsf{c}}}a_{ij}.$ Arguably the simplest hill-climbing procedure would start from some partition, perhaps a random one, and then iteratively swap the pair of vertices, one in each subset of the current partition, that makes the cut decrease the most, and stop when no swap makes the cut decrease. Carson and Impagliazzo [6] propose such an algorithm, which although relatively simple, is shown to be optimal for partitioning a planted bisection up to a poly-logarithmic factor. The procedure involves a careful initialization so that the hill-climbing part, meaning the swapping of pairs, may be seen as a refinement. This is also the case in other algorithms such as those of [17, 9, 23]. Sometimes the swapping is randomized [15, 16, 11]. Hill-climbing algorithms for graph partitioning date back at least to the work of Kernighan and Lin [19], who propose a more sophisticated swapping protocol than the one described above. The partition around medoids (PAM) algorithm of Kaufman and Rousseeuw [18] includes a hill-climbing component.
@@
In the present paper, we attach a totally different meaning to ‘hill-climbing’ that comes from the problem of clustering points in a Euclidean space. We adopt the definition of clustering proposed by Fukunaga and Hostetler [12]. (We note that this definition has been shown in [3, 4] to be intimately related to the level set definition of clustering proposed by Hartigan [14].) Assuming the points were generated iid from some density which is regular enough for the associated gradient ascent flow to be well-defined, the space, in fact the density support, is partitioned according to the basins of attraction of the density. The hill-climbing characterization comes from being a gradient method: we move each point in the density support upward, in the landscape given by the density, following the steepest ascent, and points that end at the same location are grouped together. Of course, in practice the density is unknown and one needs to find workarounds. The most famous one is the mean-shift algorithm [12, 8], although a number of other proposals have been suggested, including the most straightforward one which consists in first estimating the density, e.g., by kernel density estimation, and then computing the basin of attraction of the estimated density using a Euler scheme or similar — see [2], where a number of these procedures are analyzed.
@@
The bridge between the discrete setting of a graph and the continuous setting of a Euclidean space is the usual one: geometric graphs, meaning, graphs whose node set is embedded in a Euclidean space. Our focus will be random geometric graphs with nodes corresponding to points that are generated iid from a density with sufficient regularity, and edges connecting all pairs $\{y_{i},y_{j}\}$ with $||y_{j}-y_{j}||\leq\epsilon$. The basic question that motivates the present work is the following: Given a random geometric graph, is it possible to partition the graph in a way which is consistent with the clustering given by gradient ascent flow of the density?
@@
We answer this question affirmatively in the following sense: We propose an algorithm for graph partitioning which, when applied to a random geometric graph with underlying points generated iid from a density with enough regularity, the output of the algorithm is essentially consistent with the partition given by the basins of attraction of the gradient ascent flow of the density in an asymptotic setting, driven by $n\to\infty$, where the connectivity radius of the graph tends to zero slowly enough. Importantly, the algorithm we propose does not attempt to embed the graph. Although a strategy that would consist in embedding the graph and then applying a method for gradient ascent flow clustering is tempting, the problem of embedding the graph is arguably harder than the partitioning problem that we consider here.
@@
For more connections to the literature, we can single out the algorithm of Koontz et al. [21], which although proposed to cluster points in space, at its core is a hill-climbing graph partition algorithm close in intention to ours — but as pointed out in [2], where it is called Max Slope Shift, the algorithm in its original form fails to achieve its purpose. Our algorithm can also be seen as a special case of a class of algorithms proposed by Strazzeri and Sánchez-García [26], although the intention there is very different. We elaborate on these works later on in Section 3.2.
@@ 2 Preliminaries
In the gradient flow approach to clustering, which originated in the work of Fukunaga and Hostetler [12], each $x\in{\rm supp}(f)$ is assigned to the limit point following along the direction of the gradient of $f$. Formally, assume that $\nabla f$ is Lipschitz so that the gradient ascent flow exists. Then the gradient flow line originating at a point $x$ is the curve $\gamma_{x}$ defined by
$\gamma_{x}(0)=x;\qquad\dot{\gamma}_{x}(t)=\nabla f(\gamma_{x}(t)),\quad t\geq 0.$ (2.1)
@@
Then basin of attraction of a point $x^{*}$ is $\{x:\gamma_{x}(\infty)=x^{*}\}$. Note that the basin of attraction of $x^{*}$ is empty unless $x^{*}$ is a critical point. It turns out that, if $f$ is a Morse function [22], meaning that it is twice differentiable and its Hessian is non-singular at every critical point, then the basins of attraction corresponding to local maxima, sometimes called modes, provide a partition of the support of $f$ up to a set of zero measure.
@@
A number of mode-seeking algorithms have been proposed to approximate the gradient flow lines [5], some of them analyzed and proved to be consistent in [2]. In the same paper, another algorithm was proposed, called Max Shift, which consists in iteratively moving to the point with highest (estimated) density value. With a search radius $r>0$, which is the only tuning parameter of the method, starting at a given point a sequence is constructed by iteratively including a point within distance $r$ with highest density value: if the starting point is $x_{0}$, then Max Shift computes the sequence $(x_{k})$ given by
$x_{k+1}\in\operatorname*{arg\,max}\limits_{x\in\bar{B}(x_{k},r)}f(x),\quad k\geq 0,$ (2.2)
@@
where $\bar{B}(y,r)$ is the closed ball centered at $y$ of radius $r$. Ties are broken in some arbitrary but deterministic way. The algorithm is hill-climbing in the sense that at each step it ‘climbs’ the landscape given by the density. Max Shift clusters together points whose sequence have the same endpoint.
@@
In practice, the density $f$ is unknown and the straightforward strategy is to apply the procedure to an estimate of the density. In detail, given data points $\mathcal{Y}=\{y_{1},...,y_{n}\}$ assumed to be iid from $f$, the Max Shift method utilizes a kernel density estimator $\hat{f}_{\epsilon}$ with bandwidth $\epsilon$ to estimate the density and proceeds as before but with $\hat{f}_{\epsilon}$ in place of $f$. In the so-called medoid variant, only data points are considered, resulting in the following procedure: starting at some data point $y_{0}\in\mathcal{Y}$,
$y_{k+1}\in\operatorname*{arg\,max}\limits_{y\in\mathcal{Y}\cap\bar{B}(y_{k},r)}\hat{f}_{\epsilon}(y),\quad k\geq 0.$ (2.3)
@@
Max Shift and its medoid variant are shown in [2] to be asymptotically consistent — in the sense of Fukunaga and Hostetler [12] — under some conditions on the underlying density $f$ and on the kernel and bandwidth defining the density estimate. Specifically, it is assumed the density converges to zero at infinity; that it is twice differentiable with uniformly continuous zeroth, first, and second derivatives; and that it is of Morse regularity. And it is assumed that the kernel used is of second order and the bandwidth is chosen so that the density estimator is second order consistent, meaning that $\epsilon=\epsilon_{n}\to 0$ in such a way that, as $n\to\infty$,
$\begin{gathered}\sup_{x}|\hat{f}_{\epsilon}(x)-f(x)|\to 0,\qquad\sup_{x}\|\nabla\hat{f}_{\epsilon}(x)-\nabla f(x)\|\to 0,\qquad\sup_{x}\|\nabla^{2}\hat{f}_{\epsilon}(x)-\nabla^{2}f(x)\|\to 0.\end{gathered}$ (2.4)
@@
Our proposed algorithm for graph clustering, introduced later in Section 3, coincides with Max Shift (based on a particular kernel density estimator) in the case of a geometric graph. However, the consistency of Max Shift established in [2] do not extend to the present setting, as the kernel function implicitly used does not satisfy the required conditions. We make this point again in Remark 3.2.
@@ 3.1 Description
We describe the graph clustering method for a general graph with $n$ nodes, taken to be $[n]:=\{1,\dots,n\}$ without loss of generality, and adjacency matrix $A=(a_{ij})$ where, as usual, $a_{ij}=1$ exactly when $i$ and $j$ are neighbors in the graph, and $=0$ otherwise. We will use the convention that $a_{ii}=1$ for all $i$, so that each node is a neighbor to itself. We will sometimes use the notation $i\leftrightarrow j$ to indicate that $i$ and $j$ are neighbors, i.e., $a_{ij}=1$. The graph could be, but is not necessarily a geometric graph. The degree function is defined as usual, $q_{i}=\#\{j:j\leftrightarrow i\}$.
@@
$i_{0}=i;\qquad i_{t}\in\operatorname*{arg\,max}\{q_{j}:j\leftrightarrow i_{t-1}\},\quad t\geq 1.$ (3.1)
Ties are broken in some arbitrary, but deterministic way. We call the path $(i_{t})$ the hill-climbing path originating from $i$ where the landscape is given by the degree function $q$. In plain words, Graph Max Shift iteratively moves to neighbor with highest the degree. We then cluster together nodes whose hill-climbing paths end at the same node. Finally, we merge together any two clusters whose associated nodes are within $\tau$ hops, where $\tau\geq 1$ is the only tuning parameter of the method.
@@ 3.2.1 Connection to Max Shift
Our claim is that Graph Max Shift implements gradient ascent flow clustering when applied to a geometric graph. Let $\mathcal{Y}:=\{y_{1},\dots,y_{n}\}$ be a set of points in some Euclidean space, say, $\mathbb{R}^{d}$, and consider the neighborhood graph with connectivity radius $\epsilon$, denoted $\mathcal{G}(\mathcal{Y};\epsilon)$, defined by the adjacency matrix $a_{ij}=1$ if $\|y_{i}-y_{j}\|\leq\epsilon$, and $=0$ otherwise.
@@
We remark that, in the situation where the user is only provided with an adjacency matrix, which is the one we envision, when applying Graph Max Shift the user does not have a choice of kernel function or bandwidth.
@@ 3.2.2 Connection to Max Slope Shift
Koontz et al. [21] proposed an algorithm for clustering points in space, which is called Max Slope Shift in [2] for reasons that will become clear shortly. Given a point set $\mathcal{Y}:=\{y_{1},\dots,y_{n}\}$, it forms the weighted neighborhood graph with connectivity radius $\epsilon$, namely, the graph on $[n]$ with dissimilarity between nodes $i,j\in[n]$ given by $\delta_{ij}=\|y_{i}-y_{j}\|$ if $\|y_{i}-y_{j}\|\leq\epsilon$, and $\delta_{ij}=\infty$ otherwise. Here $i$ and $j$ are considered neighbors if $\delta_{ij}<\infty$, and the degree function is defined as before, $q_{i}=\#\{j:j\leftrightarrow i\}$. Then, starting at node $i$, Max Slope Shift computes the following sequence of nodes:
$i_{0}=i;\qquad i_{t}\in\operatorname*{arg\,max}\left\{\frac{q_{j}-q_{i}}{\|y_{i}-y_{j}\|}:j\leftrightarrow i_{t-1}\right\},\quad t\geq 1.$ (3.4)
As argued in [2], it is not hard to see that the end points will be local maxima not of the density itself, but of its gradient, meaning, density inflection points. The algorithm thus fails to output an accurate clustering as defined by Fukunaga and Hostetler [12]. (A regularization is proposed in [2] to remedy this behavior.)
Max Slope Shift can be nonetheless seen to be closely related to ours. For one, it implies the use of the flat kernel for density estimation; and two, if one were to imagine how it would be adapted to clustering an unweighted graph, one would end up with Graph Max Shift, as the natural thing to do in that case is to replace the Euclidean distance in the denominator in the fraction appearing in (3.4) with 1.

### v2

@@ Abstract
We present a method for graph clustering that is analogous to gradient ascent methods previously proposed for clustering points in space. The algorithm, which can be viewed as a max-degree hill-climbing procedure on the graph, iteratively moves each node to a neighboring node of highest degree. We show that, when applied to a random geometric graph whose nodes correspond to data drawn i.i.d. from a density with Morse regularity, the method is asymptotically consistent. Here, consistency is in the sense of Fukunaga and Hostetler, meaning, with respect to the partition of the support of the density defined by the basins of attraction of the density gradient flow.
@@ 1 Introduction
A hill-climbing algorithm is typically understood as an algorithm that makes ‘local’ moves. In a sense, this class of procedures is the discrete analog of the class of gradient-based and higher-order methods in continuous optimization. Such algorithms have been proposed in the context of graph partitioning, sometimes as a refinement step, where the objective function is typically a notion of cut and local moves often take the form of swapping vertices in order to improve the value of the objective function. More specifically, consider an undirected graph consisting of $n$ nodes, which we take to be $n]:=\{1,\dots,n\}$ without loss of generality, and adjacency matrix $A=(a_{ij})$, so that $a_{ij}=1$ when $i$ and $j$ are connected in the graph, and $a_{ij}=0$ otherwise. Consider the simplest variant of graph partitioning, where assuming $n$ is even, the goal is to bisect the node set into two subsets, $I\subset[n]$ and $I^{\mathsf{c}}$, in order to minimize the cut: ${\rm cut}(I)=\sum_{i\in I}\sum_{j\in I^{\mathsf{c}}}a_{ij}.$ Arguably the simplest hill-climbing procedure would start from some partition, perhaps a random one, and then iteratively swap the pair of vertices, one in each subset of the current partition, that makes the cut decrease the most, and stop when no swap makes the cut decrease. Carson and Impagliazzo (2001) propose such an algorithm, which although relatively simple, is shown to be optimal for partitioning a planted bisection in the sense of achieving the information theoretic recovery threshold up to a poly-logarithmic factor. The procedure involves a careful initialization so that the hill-climbing part, meaning the swapping of pairs, may be seen as a refinement. This is also the case in other algorithms such as those of Karypis and Kumar (1998); Condon and Karp (1999); Mossel et al. (2015). Sometimes the swapping is randomized Jerrum and Sorkin (1993); Juels (1996); Dimitriou and Impagliazzo (1998). Hill-climbing algorithms for graph partitioning date back at least to the work of Kernighan and Lin (1970), who propose a more sophisticated swapping protocol than the one described above. The partition around medoids (PAM) algorithm of Kaufman and Rousseeuw (1987) includes a hill-climbing component.
@@
In the present paper, we attach a totally different meaning to ‘hill-climbing’ that comes from the problem of clustering points in a Euclidean space as proposed by Fukunaga and Hostetler (1975). (We note that this definition has been shown in Arias-Castro and Qiao (2023b, a) to be intimately related to the level set definition of clustering proposed by Hartigan (1975).) Assuming the points were generated iid from some density which is regular enough for the associated gradient ascent flow to be well-defined, the space, in fact the density support, is partitioned according to the basins of attraction of that gradient flow. The hill-climbing characterization comes from being a gradient method: we move each point in the density support upward, in the landscape given by the density, following the steepest ascent, and points that end at the same location are grouped together. Of course, in practice the density is unknown and one needs to find workarounds. The most famous one is the mean-shift algorithm Fukunaga and Hostetler (1975); Cheng (1995), although a number of other proposals have been suggested, including the most straightforward one which consists in first estimating the density, e.g., by kernel density estimation, and then computing the basin of attraction of the estimated density using a Euler scheme or similar — see Arias-Castro and Qiao (2025), where a number of these procedures are analyzed.
@@
The bridge between the discrete setting of a graph and the continuous setting of a Euclidean space is the usual one: geometric graphs, meaning, graphs whose node set is embedded in a Euclidean space. Our focus will be random geometric graphs with nodes corresponding to points that are generated iid from a density with sufficient regularity, and edges connecting all pairs of nodes whose associated points are within distance $\epsilon$. The basic question that motivates the present work is the following:
Given the adjacency matrix of a random geometric graph, is it possible to partition the graph in a way which is consistent with the clustering given by the gradient ascent flow of the density?
We emphasize that the geometric graph is only known through its adjacency matrix. No embedding of the graph is provided, in other words, the points associated with the nodes are unknown. (In this context, these points are often referred to as ‘latent positions’.)
@@
We answer this question affirmatively in the following sense: We propose an algorithm for graph partitioning which, when applied to a random geometric graph with underlying points generated iid from a density with enough regularity, outputs a partition which is consistent with the partition given by the basins of attraction of the gradient ascent flow of the density in an asymptotic setting, driven by $n\to\infty$, where the connectivity radius of the graph tends to zero slowly enough. Importantly, the algorithm we propose does not attempt to embed the graph. Although a strategy that would consist in embedding the graph and then applying a method for gradient ascent flow clustering is tempting, the problem of embedding the graph is arguably harder than the partitioning problem that we consider here.
@@
For more connections to the literature, we can single out the algorithm of Koontz et al. (1976), which although proposed to cluster points in space, at its core is a hill-climbing graph partition algorithm close in intention to ours — but as pointed out in Arias-Castro and Qiao (2025), where it is called Max Slope Shift, the algorithm in its original form fails to achieve its purpose. Our algorithm can also be seen as a special case of a class of algorithms proposed by Strazzeri and Sánchez-García (2022), although the intention there is very different. We elaborate on these works later on in Section 3.2.
@@ 2 Preliminaries
In the gradient flow approach to clustering, which originated in the work of Fukunaga and Hostetler (1975), each $x\in{\rm supp}(f)$ is assigned to the limit point following along the direction of the gradient of $f$. Formally, assume that $\nabla f$ is Lipschitz so that the gradient ascent flow exists. Then the gradient flow line originating at a point $x$ is the curve $\gamma_{x}$ defined by
$\gamma_{x}(0)=x;\hskip 18.49988pt\dot{\gamma}_{x}(t)=\nabla f(\gamma_{x}(t)),\qquad t\geq 0.$ (2.1)
@@
The basin of attraction of a point $x^{*}$ is $\{x:\gamma_{x}(\infty)=x^{*}\}$. Note that the basin of attraction of $x^{*}$ is empty unless $x^{*}$ is a critical point. It turns out that, if $f$ is a Morse function (Milnor, 1963), meaning that it is twice differentiable and its Hessian is non-singular at every critical point, then the basins of attraction corresponding to local maxima, sometimes called modes, provide a partition of the support of $f$ up to a set of zero measure. See, e.g., (Arias-Castro and Qiao, 2025, Lem 2.2), which is argued entirely based on (Banyaga and Hurtubise, 2013, Cor 3.3, Th 4.2).
@@
A number of mode-seeking algorithms have been proposed to approximate the gradient flow lines Carreira-Perpiñán (2015), some of them analyzed and proved to be consistent in Arias-Castro and Qiao (2025). In the same paper, another algorithm was proposed, called Max Shift, which consists in iteratively moving to the point with highest (estimated) density value. With a search radius $r>0$, which is the only tuning parameter of the method, starting at a given point a sequence is constructed by iteratively including a point within distance $r$ with highest density value: if the starting point is $x_{0}$, then Max Shift computes the sequence $(x_{k})$ given by
$x_{k+1}\in\operatorname*{arg\,max}\big\{f(x):x\in\bar{B}(x_{k},r)\big\},\qquad k\geq 0,$ (2.2)
@@
where $\bar{B}(y,r)$ is the closed ball centered at $y$ of radius $r$. Ties are broken in some arbitrary but deterministic way, e.g., by following the lexicographic order. The algorithm is hill-climbing in the sense that at each step it ‘climbs’ the landscape given by the density. Max Shift clusters together points whose sequence have the same endpoint.
@@
In practice, the density $f$ is unknown and the straightforward strategy is to apply the procedure to an estimate of the density. In detail, given data points $\mathcal{Y}=\{y_{1},\dots,y_{n}\}$ assumed to be iid from $f$, the Max Shift method utilizes a kernel density estimator $\hat{f}_{\epsilon}$ with bandwidth $\epsilon$ to estimate the density and proceeds as before but with $\hat{f}_{\epsilon}$ in place of $f$. In the so-called medoid variant, only data points are considered, resulting in the following procedure: starting at some data point $y_{i_{0}}\in\mathcal{Y}$,
$y_{i_{k+1}}\in\operatorname*{arg\,max}\big\{\hat{f}_{\epsilon}(y):y\in\mathcal{Y}\cap\bar{B}(y_{i_{k}},r)\big\},\qquad k\geq 0.$ (2.3)
For references on medoid variants of various hill-climbing clustering methods, see (Arias-Castro and Qiao, 2025, Appendix B).
@@
Max Shift and its medoid variant are shown in Arias-Castro and Qiao (2025) to be asymptotically consistent in the sense of Fukunaga and Hostetler (1975) (formally defined in Section 3.3) under some conditions on the underlying density $f$ and on the kernel and bandwidth defining the density estimate. Specifically, it is assumed the density converges to zero at infinity; that it is twice differentiable with uniformly continuous zeroth, first, and second derivatives; and that it is of Morse regularity; and it is assumed that the kernel used is of second order and the bandwidth is chosen so that the density estimator is second order consistent, meaning that $\epsilon=\epsilon_{n}\to 0$ in such a way that, as $n\to\infty$,
$\begin{gathered}\sup_{x}|\hat{f}_{\epsilon}(x)-f(x)|\to 0,\hskip 18.49988pt\sup_{x}\|\nabla\hat{f}_{\epsilon}(x)-\nabla f(x)\|\to 0,\hskip 18.49988pt\sup_{x}\|\nabla^{2}\hat{f}_{\epsilon}(x)-\nabla^{2}f(x)\|\to 0.\end{gathered}$ (2.4)
@@
Our proposed algorithm for graph clustering, introduced later in Section 3, coincides with Max Shift (based on a particular kernel density estimator) in the case of a geometric graph. However, the consistency of Max Shift established in Arias-Castro and Qiao (2025) does not extend to the present setting, as the kernel function that Max Shift implicitly uses does not satisfy the conditions required Arias-Castro and Qiao (2025).
@@ 3.1 Description
We describe the graph clustering method for a general graph with $n$ nodes, taken to be $n]:=\{1,\dots,n\}$ without loss of generality, and adjacency matrix $A=(a_{ij})$ where, as usual, $a_{ij}=1$ exactly when $i$ and $j$ are neighbors in the graph, and =0=0 otherwise. We will use the convention that $a_{ii}=1$ for all $i$, so that each node is a neighbor to itself. We will sometimes use the notation $i\leftrightarrow j$ to indicate that $i$ and $j$ are neighbors, i.e., $a_{ij}=1$. The graph could be, but is not necessarily a geometric graph. The degree function is defined as usual, $q_{i}=\#\{j:j\leftrightarrow i\}$.
@@
$i_{0}=i;\hskip 18.49988pti_{t}\in\operatorname*{arg\,max}\{q_{j}:j\leftrightarrow i_{t-1}\},\qquad t\geq 1.$ (3.1)
Ties are broken in some arbitrary, but deterministic way, e.g., by choosing the node with highest numerical index among maximizers. We call the path $(i_{t})$ the hill-climbing path originating from $i$ where the landscape is given by the degree function $q$. In plain words, Graph Max Shift iteratively moves to a neighbor with highest degree. We then cluster together nodes whose hill-climbing paths end at the same terminal node. Optionally, we merge together any two clusters whose associated terminal nodes are within $\tau$ hops — in which case $\tau\geq 1$ is a tuning parameter of the method (the only one).
@@ 3.2.1 Connection to Max Shift
Our claim is that Graph Max Shift implements gradient ascent flow clustering when applied to a geometric graph. Let $\mathcal{Y}:=\{y_{1},\dots,y_{n}\}$ be a set of points in some Euclidean space, say, $\mathbb{R}^{d}$, and consider the neighborhood graph with connectivity radius $\epsilon$, denoted $\mathcal{G}(\mathcal{Y};\epsilon)$, defined by the adjacency matrix $a_{ij}=1$ if $\|y_{i}-y_{j}\|\leq\epsilon$, and =0=0 otherwise.
@@
When started at $y^{0}\in\mathcal{Y}$, the sequence built by Graph Max Shift, from the point of view of the latent positions, is given by
$y^{k+1}\in\operatorname*{arg\,max}\limits_{y\in\mathcal{Y}\cap\bar{B}(y^{k},\epsilon)}\hat{f}_{\epsilon}(y),\qquad k\geq 0,$ (3.4)
which is exactly the sequence computed by Max Shift applied to $\mathcal{Y}$ based on the kernel density estimator $\hat{f}_{\epsilon}$. (Note that $y^{k}$ above corresponds to $y_{i_{k}}$ for some $i_{k}\in[n]$.)
@@
We remark that, in the situation where the user is only provided with an adjacency matrix, which is the one we envision, when applying Graph Max Shift the user does not have a choice of kernel function $K$ or bandwidth $\epsilon$.
@@ 3.2.2 Connection to Max Slope Shift
Koontz et al. (1976) proposed an algorithm for clustering points in space, which is called Max Slope Shift in Arias-Castro and Qiao (2025) for reasons that will become clear shortly. Given a point set $\mathcal{Y}:=\{y_{1},\dots,y_{n}\}$, it forms the weighted neighborhood graph with connectivity radius $\epsilon$, namely, the graph on $[n]$ with dissimilarity between nodes $i,j\in[n]$ given by $\delta_{ij}=\|y_{i}-y_{j}\|$ if $\|y_{i}-y_{j}\|\leq\epsilon$, and $\delta_{ij}=\infty$ otherwise. Here $i$ and $j$ are considered neighbors if $\delta_{ij}<\infty$, and the degree function is defined as before, $q_{i}=\#\{j:j\leftrightarrow i\}$. Then, starting at node $i$, Max Slope Shift computes the following sequence of nodes:
$i_{0}=i;\hskip 18.49988pti_{t}\in\operatorname*{arg\,max}\left\{\frac{q_{j}-q_{i_{t-1}}}{\|y_{j}-y_{i_{t-1}}\|}:j\leftrightarrow i_{t-1}\right\},\qquad t\geq 1.$ (3.5)
As argued in Arias-Castro and Qiao (2025), it is not hard to see that the end points will be local maxima not of the density itself, but of its gradient, meaning, density inflection points. The algorithm thus fails to output an accurate clustering as defined by Fukunaga and Hostetler (1975). (A regularization is proposed in Arias-Castro and Qiao (2025) to remedy this behavior.)
Max Slope Shift can be nonetheless seen to be closely related to ours. Indeed, it implies the use of the flat kernel for density estimation; and, if one were to imagine how it would be adapted to clustering an unweighted graph, one would end up with Graph Max Shift, as the natural thing to do in that case is to replace the Euclidean distance in the denominator in the fraction appearing in (3.5) with 1.